# ConKeD: Multiview contrastive descriptor learning for keypoint-based retinal image registration

### v1

@@ 3.3 Comparison with state-of-the-art registration methods
FIRE A P S Avg. W. Avg.
VOTUS Motta et al. (2019) 0.812 0.681 0.672 0.934 0.762 0.811
SR Manual Liu et al. (2022) - 0.783 0.542* 0.94 0.755* 0.780*
SR PBO Liu et al. (2022) - 0.789 0.516* 0.944 0.750* 0.773*
Proposed 0.758 0.749 0.489 0.945 0.728 0.758
REMPE Hernandez-Matas et al. (2020) 0.773 0.66 0.542 0.958 0.72 0.774
Retina-R2D2 Rivas-Villar et al. 0.695 0.726 0.352 0.925 0.645 0.575
Rivas-Villar Rivas-Villar et al. (2022) 0.657 0.660 0.293 0.908 0.620 0.552
@@
Table 5: Comparison between our approach and state of the art methods, sorted by average. * indicates that the results were calculated using one image less. Best results in bold. Results sourced from the referenced papers.
@@
The comparison with the state-of-the-art is depicted in Table 5. In this comparison, we include the three best-performing approaches on the FIRE dataset as well as previous approaches using deep learning. The results show that the best performing method overall is VOTUS Motta et al. (2019) and the best deep learning method is SuperRetina Liu et al. (2022). However, it should be noted that SuperRetina does not provide the AUC for the whole FIRE dataset as it is done in every other work in the state of the art Motta et al. (2019); Hernandez-Matas et al. (2020); Rivas-Villar et al. (2022). Analyzing each category separately, we can see that in A our proposal obtains the second best results only behind SuperRetina by around 4%. In category P both our method and SuperRetina produce significantly worse results than VOTUS Motta et al. (2019), a classical method that uses a quadratic transformation to help improving the alignment in these low overalapping cases. Finally, in category S, which is the easiest to register, the results of most methods are within a margin of around 1%, therefore the differences are not notable.
@@
In terms of Registration Score, our method is the second best deep learning approach and overall provides results that are close to those reported by SuperRetina. However, our method provides several advantages and there are also some important experimental differences between both works that should be considered. SuperRetina is trained using a private dataset with more than double the images than the public DRIVE dataset used in our work. Furthermore, their approach requires extensive ad-hoc image pre-processing. Moreover, the descriptor computed by SuperRetina has 256 channels compared to 128 in our descriptor. Additionally, the results of SuperRetina depend greatly on the number of detected keypoints. SuperRetina detects, approximately, five times more keypoints than our approach. In that regard, the authors of SuperRetina also tested a variant of their method that detects approximately the same amount as our work, obtaining a notable performance drop, from 0.755 to 0.685 (Avg.) in which case our method, which produces approximately the same number of keypoints, would obtain better results. Using more keypoints usually increases performance as it provides more chances for the descriptors to match and more paired keypoints to compute a transformation, at the cost of more computation. This is specially notable in category P where the small overalapping limits the number of keypoints usable for registration. Additionally, SuperRetina requires double inference (i.e. double run), including double detection, description, descriptor matching and transformation calculation. In contrast, our method does not require any of these extra steps. Furthermore, our experiments revealed that running our method a second time does not improve the results. This evidences that our descriptors and keypoints are more robust to transformations since they are correctly calculated in a single pass, despite the larger transformations.
Regarding the different categories, the most relevant difference is in category P where our method obtains 2.7% less and 5.3% less AUC than both SuperRetina methods. As explained, this is due to the lower number of keypoints detected per image. In the other categories, where the amount of keypoints is less relevant, our results are even closer. Particularly, our results are slightly better in category S.
@@ 4 Conclusion
In order to validate our method, we perform several experiments in the well known FIRE dataset and analyze several factors in the methodology. The results show that our proposal clearly outperforms triplet loss, the alternative typically used to learn descriptors in the state of the art. Moreover, we also demonstrate the performance improvement of using multiple positives. In terms of the losses, we found that SupCon Loss and a multi-positive InfoNCE performed similarly with a slight advantage to the latter evidencing that our framework is robust to specific loss terms. Moreover, our method provides results competitive with the best deep learning state of the art method while using less training samples, less keypoints and not requiring double inference nor pre-processing.
As future work, it would be interesting to test more losses in combination with our ConKeD framework. Moreover, developing specific losses for the multi-positive paradigm is also desirable to try and improve the results. In order to improve the whole pipeline, increasing the number of keypoints while keeping their high specificity could potentially improve the results, specially in category P, which has the lowest overlapping.
@@ Acknowledgments
This work is supported by Ministerio de Ciencia e Innovación, Government of Spain, through the RTI2018-095894-B-I00, PID2019-108435RB-I00, TED2021-131201B-I00, and PDC2022-133132-I00 research projects; Consellería de Cultura, Educación e Universidade, Xunta de Galicia, through Grupos de Referencia Competitiva ref. ED431C 2020/24, predoctoral fellowship ref. ED481A 2021/147 and the postdoctoral fellowship ref. ED481B-2022-025. CITIC, Centro de Investigación de Galicia ref. ED431G 2019/01, receives financial support from Consellería de Cultura, Educación e Universidade, Xunta de Galicia, through the ERDF of the European Union (80%) and Secretaría Xeral de Universidades (20%).

### v2

@@ 3.3 Comparison with state-of-the-art registration methods
Figure 6: Representative registration examples from the FIRE dataset for our proposals and different state-of-the-art methods.
FIRE A P S Avg. W. Avg. Keypoints
Training
images
VOTUS [21] 0.812 0.681 0.672 0.934 0.762 0.811 N/A N/A
SR Manual [24] - 0.783 0.542* 0.94 0.755* 0.780* 740 97 + 844
SR PBO [24] - 0.789 0.516* 0.944 0.750* 0.773* 740 97 + 844
Proposed 0.758 0.749 0.489 0.945 0.728 0.758 115 15 & 40
REMPE [20] 0.773 0.66 0.542 0.958 0.72 0.774 - N/A
Retina-R2D2 [8] 0.695 0.726 0.352 0.925 0.645 0.575 5000 1748
Rivas-Villar [23] 0.657 0.660 0.293 0.908 0.620 0.552 115 20
@@
Table 5: Comparison between our approach and state of the art methods, sorted by average. * indicates that the results were calculated using one image less. N/A indicates that the method does not use keypoints or does not require training samples. - indicates that the method is not publicly available and does not specify the number of keypoints used. $k+d$ denotes that k images are used for the detector training and $k+d$ images are used for the descriptor training. Meanwhile, $k\&d$ denotes that $k$ images are used for the detector training and $d$ images are used for the descriptor training. Best results in bold. Results sourced from the referenced papers.
@@
The comparison with the state-of-the-art is depicted in Table 5. In this comparison, we include the three best-performing approaches on the FIRE dataset as well as previous approaches using deep learning. The results show that the best performing method overall is VOTUS [21] and the best deep learning method is SuperRetina [24]. However, it should be noted that SuperRetina does not provide the AUC for the whole FIRE dataset as it is done in every other work in the state of the art [21, 20, 23]. Analyzing each category separately, we can see that in A our proposal obtains the second best results only behind SuperRetina by around 4%. In category P both our method and SuperRetina produce significantly worse results than VOTUS [21], a classical method that uses a quadratic transformation to help improving the alignment in these low overlapping cases. Finally, in category S, which is the easiest to register, the results of most methods are within a margin of around 1%, therefore the differences are not notable.
@@
In terms of Registration Score, our method is the second best deep learning approach and overall provides results that are close to those reported by SuperRetina. Regarding the different categories, the most relevant difference is in category P where our method obtains 2.7% less and 5.3% less AUC than both SuperRetina methods. As explained, this is due to the lower number of keypoints detected per image. In the other categories, where the amount of keypoints is less relevant, our results are even closer, even surpassing SuperRetina on in category S. Importantly, our method provides several advantages and there are also some important experimental differences between both works that should be considered. SuperRetina trains its detector using a private dataset containing 97 images whereas we train the detector with just 15. Additionally, SuperRetina uses an extra set of 844 images to train the descriptor, for a total of 941, whereas we use just 40 images for our descriptor. Furthermore, SuperRetina requires extensive ad-hoc image pre-processing. Moreover, the descriptor computed by SuperRetina has 256 channels compared to 128 in our descriptor. Additionally, the results of SuperRetina depend greatly on the number of detected keypoints. As shown on Table 5, SuperRetina detects more than 6×6$\times$6 × more keypoints than our approach. In that regard, the authors of SuperRetina also tested a variant of their method that detects approximately the same amount as our work (using just the labeled keypoints in their training set), obtaining a notable performance drop, from 0.755 to 0.685 (Avg.) in which case our method, which produces approximately the same number of keypoints, would obtain better results. Using more keypoints usually increases performance as it provides more chances for the descriptors to match and more paired keypoints to compute a transformation, at the cost of more computation. This is especially notable in category P where the small overlapping limits the number of keypoints usable for registration. Additionally, SuperRetina requires double inference (i.e. double run), including double detection, description, descriptor matching and transformation calculation. In contrast, our method does not require any of these extra steps making it more efficient.
Our method was designed with efficiency in mind as our multi-positive multi-negative training approach allows us to leverage more information during training, allowing for increased data efficiency. Additionally, through the use of domain-specific keypoints we can limit the cost of matching and transformation estimation. Overall, our proposed approach demonstrates execution times that are suitable for use in clinical day-to-day practice, providing fast image registrations. In particular, the average execution time for our whole methodology, including each step in the process is, approximately, 0.088 seconds per image pair. This time can be itself divided into the different steps of the method: 0.03 seconds for keypoint detection inference, 0.01 seconds for descriptor inference, 0.0006 for descriptor matching and 0.047 for transformation computation through RANSAC. In relation to this, our method detects, on average, 115 keypoints throughout the FIRE dataset. Given that the complexity of the descriptor matching stage is $O(n^{2})$, this means that $13225115^{2}=$ 13225 comparisons are performed for each image pair. However, if our approach required the same amount of keypoints as SuperRetina (740 keypoints) or Retina-R2D2 (5000 keypoints), the number of comparisons required for the matching process would increase to 547,600 or 25,000,000, respectively. This would increase the matching time by factors of 41x and 1890x. Moreover, SuperRetina uses heavier descriptors and double inference, which would increase execution time and computational cost further. Our experiments revealed that running our method a second time does not improve the results. This evidences that our descriptors and keypoints are more robust to transformations since they are correctly calculated in a single pass, despite the larger transformations.
Finally, representative examples of registered images from the FIRE dataset using our approach as well as available state-of-the-art methods can be seen in Figure 6. While there are slight numerical differences between our approach and others, such as SuperRetina, they can hardly be visually spotted in the images. This is due to the high resolution of the images in FIRE. Small registration errors (i.e. 1 or 2 pixels) become really hard to spot on images of such size while significantly impacting numerical results.
@@ 4 Conclusion
In order to validate our method, we perform several experiments in the well-known FIRE dataset and analyze several factors in the methodology. The results show that our proposal clearly outperforms triplet loss, the alternative typically used to learn descriptors in the state of the art. In terms of the losses, we found that SupCon Loss and a multi-positive InfoNCE performed similarly with a slight advantage to the latter evidencing that our framework is robust to specific loss terms. Moreover, our method provides results competitive with the best deep learning state of the art method while using less training samples, less keypoints and not requiring double inference nor pre-processing.
As future work, it would be interesting to test more losses in combination with our ConKeD framework. In this regard, developing specific losses for the multi-positive paradigm is also desirable to try and improve the results. Additionally, a limitation of our current approach and of others in the state of the art is the reliance on inherently limited domain-specific keypoints, such as crossovers and bifurcations. These keypoints are not distributed evenly across the retina and may be completely absent in some parts. This can reduce the overall registration performance since there are no points to accurately drive the transformation in every part of the fundus. The arbitrary nature of these domain-specific keypoints may also reduce the performance of these methods on patients with less crossovers and bifurcations visible in the eye fundus. Therefore, future works could focus on developing methods to increase the number of keypoints while keeping a high specificity. This could potentially improve the results, especially in category P, which has the lowest overlapping. In that regard, it could even be convenient to study the inclusion of generic keypoints. These, despite their generally lower specificity, could help regularize the registration if there are zones of the image that lack domain-specific keypoints. Therefore, the inclusion of generic keypoints in a measured way may be beneficial without a significant reduction in the efficiency of the method.
Declarations
The authors have no relevant financial or non-financial interests to disclose.
@@ Acknowledgments
This work is supported by Ministerio de Ciencia e Innovación, Government of Spain, through the PID2019-108435RB-I00, TED2021-131201B-I00, and PDC2022-133132-I00 research projects; Consellería de Cultura, Educación e Universidade, Xunta de Galicia, through Grupos de Referencia Competitiva ref. ED431C 2020/24, predoctoral fellowship ref. ED481A 2021/147 and the postdoctoral fellowship ref. ED481B-2022-025. Funding for open access charge: Universidade da Coruña/CISUG.